Please help!


Expert Q&A Done 1 .) A 5% (w/v) dextrose IV fluid (D5W) contains 20 mEq K+ per liter of solution. If a patient received 420 mL of this IV solution, how many millimoles of K+ were delivered? 2.) A 25.0 mL multidisciplinar vial of a medication is available containing 500mg of a drug dissolves in normal saline. What is % (w/v) concentration of this preparation? 3.) How many grams of insulin are present in 5.0g if 0.450% (w/w) solution? Can you please show how to do this by using conversion factors/dimensional analysis? Thank you!
Answer #1

1. Potassium (K+) is a univalent ion and so 1mEq = 1 mmol

D5W contains 20mEq/ Liter K+

So 420ml of D5W contains : 420 (ml)/1000(ml) X 20 = 8.4mmol of K+

2. Concentration = weight/ Volume

Drug concentration in the given solution = 500(mg)/25(ml) = 20%

3. w/w = weight/weight

0.450% (w/w)= 0.450g/100g

Therefore 5g contains: 0.450/100x5=0.0225g of insulin